# Random Lévy Looptrees and Lévy Maps

### v2

@@ 4.1.   Hölder continuity of the distances
Recall the moment bound on the Brownian labels (2.9), then if the Hölder continuity (4.3) on the looptree distances held true with a deterministic constant $C_{a}$, then we could apply the classical Kolmogorov continuity criterion, showing that for any $a<1/\boldsymbol{\eta}$ and any $q>0$, the process $(Z^{\mathrm{ex}}_{t})_{t}$ has a modification that is Hölder continuous with any exponent $b<q^{-1}(aq/2-1)$, which is arbitrarily close to $1/(2\boldsymbol{\eta})$ when $q$ is large. The idea then is to adapt the proof of this criterion, relying on (4.4) where the constant $K_{a}$ is indeed deterministic.
As in the previous proof, it suffices to prove the claim in the infinite volume setting by replacing $X^{\mathrm{ex}}$ by the unconditioned Lévy process $X$ and constructing the associated looptree and label process $Z$. By absolute continuity, it then extends to the bridge, and finally to the excursion by the Vervaat transform. Fix $a<1/\boldsymbol{\eta}$ and $q>0$ and let $b<q^{-1}(aq/2-1)$ and let $A_{n}$ be the event
$A_{n}=\Bigl{\{}\sup_{0\leq i\leq 2^{n}-1}d_{\mathscr{L}(X)}(i2^{-n},(i+1)2^{-n})\leq K_{a}2^{-an}\Bigr{\}}$
from (4.4), where $K_{a}$ is a constant chosen such that $\sum_{n}\mathop{}\!\mathbb{P}(A_{n}^{c})<\infty$. Then using a union bound, the Markov inequality for the conditional expectation given the process $X$, and the moment bound (2.9), we have for every $n\geq 1$:
$\mathop{}\!\mathbb{P}\Bigl{(}A_{n}\text{ and }\sup_{0\leq i\leq 2^{n}-1}|Z_{i2^{-n}}-Z_{(i+1)2^{-n}}|>2^{-nb}\Bigr{)}$ $\leq 2^{n}\sup_{0\leq i\leq 2^{n}-1}\mathop{}\!\mathbb{P}(A_{n}\text{ and }|Z_{i2^{-n}}-Z_{(i+1)2^{-n}}|^{q}>2^{-nbq})$
$\leq 2^{n}\sup_{0\leq i\leq 2^{n}-1}2^{nbq}\mathop{}\!\mathbb{E}[\mathbf{1}_{\{A_{n}\}}\mathop{}\!\mathbb{E}[|Z_{i2^{-n}}-Z_{(i+1)2^{-n}}|^{q}\mid X]]$
$\leq 2^{n}\sup_{0\leq i\leq 2^{n}-1}2^{nbq}K_{q}\mathop{}\!\mathbb{E}[\mathbf{1}_{\{A_{n}\}}d_{\mathscr{L}(X)}(i2^{-n},(i+1)2^{-n})^{q/2}]$
$\leq 2^{n}2^{nbq}K_{q}K_{a}^{q/2}2^{-anq/2}$
$\leq K_{q}K_{a}^{q/2}2^{n(1+bq-aq/2)},$
which is a convergent series since $01+bq-aq/2<$ 0. Since $\sum_{n}\mathop{}\!\mathbb{P}(A_{n}^{c})<\infty$, then we infer that
$\sum_{n\geq 1}\mathop{}\!\mathbb{P}\Bigl{(}\sup_{0\leq i\leq 2^{n}-1}|Z_{i2^{-n}}-Z_{(i+1)2^{-n}}|>2^{-nb}\Bigr{)}<\infty.$
From here we can catch up with the proof of the classical Kolmogorov continuity criterion: the Borel–Cantelli lemma shows that almost surely, these events fail eventually; we then extend the bound to any pair of dyadic numbers and then to any times by density. ∎
From now on, we always deal with this Hölder continuous modification of $Z$ and $Z^{\mathrm{ex}}$. This further implies that the distances in the map are Hölder continuous.
@@
For any $a\in(0,1/(2\boldsymbol{\eta}))$, the projection $\mathbf{p}_{M}$ is almost surely Hölder continuous with exponent $a$ in the sense that almost surely, there exists $C_{a}\in(0,\infty)$ such that:
@@
As for Proposition 4.3, we shall see that $\mathbf{p}_{M}$ is almost surely not Hölder continuous with exponent larger than $1/(2\boldsymbol{\eta})$, see again Remark 4.7.
@@
Define for every $0\leq s\leq t\leq 1$ a pseudo-distance by:
$D^{\circ}(s,t)=D^{\circ}(t,s)=Z^{\mathrm{ex}}_{s}+Z^{\mathrm{ex}}_{t}-2\inf_{[s,t]}Z^{\mathrm{ex}}.$ (4.7)
Then the Hölder continuity of $Z^{\mathrm{ex}}$ from Corollary 4.5 transfers to $D^{\circ}$ which thus satisfies the upper bound (4.6). Our claim follows then from the bound
$d_{M}\leq D^{\circ}.$
Indeed, a similar bound holds in the discrete setting, see e.g. Le Gall [LG07, Lemma 3.1] in a slightly different context, and it passes to the limit. To explain it, in the construction of a discrete map from a labelled looptree, the labels on the looptree codes the distances in the map to a distinguished vertex. Precisely, the chain of successors define for every vertex a special geodesic path to the distinguished vertex. Then if one takes two arbitrary vertices and follows these two special geodesic paths starting from each of them until they meet, then the sum of the two lengths is precisely the discrete analogue of $D^{\circ}$, and is clearly not smaller than the graph distance in the map between the two vertices. ∎
@@ 5. Approximations in the continuum world
In this section we ask ourselves the following question. Suppose that for all $\lambda>0$ we are given a Lévy process $X^{(\lambda)}$ with no negative jump and infinite variation paths, with exponent $\psi^{(\lambda)}$, and another one $X$ with exponent $\psi$. Suppose that $X^{(\lambda)}$ and $X$ satisfy the integrability condition (1.3), so they all have continuous transition densities $p^{(\lambda)}_{t}(x)$ and $p_{t}(x)$ respectively, and we can define bridges and excursions as before. If these transition densities converge, then it is well known that the Lévy processes converge in law; but do the corresponding looptrees, Gaussian labels, and continuum maps converge as well? If the limit $X$ has no Gaussian part, then it has the pure jump property from [Mar22b] and Section 4 there answers positively the question. Here we consider the general setting: we prove the convergence of labelled looptrees in Section 5.1 and that of maps (along a subsequence) in Section 5.2.
A typical example of application we have in mind is the case where $X^{(\lambda)}$ is an $\alpha$-stable spectrally positive Lévy process with drift $\lambda$, with varying $\alpha$ and $\lambda$, studied in Section 6.
@@
Suppose that for every $t\geq 0$, the transition density $p^{(\lambda)}_{t}$ converges uniformly as $\lambda\to\infty$ to $p_{t}$. Then $X^{(\lambda)}\to X$ in distribution for the Skorokhod $J_{1}$ topology under the unconditional law as well as under $\mathop{}\!\mathbb{P}(\,\cdot\mid X^{(\lambda)}_{1}=\vartheta)$ for any $\vartheta\in\mathbb{R}$, and finally under the excursion law.
@@
for any continuous and bounded function $F$. We infer from the convergence in distribution of the unconditioned processes and the uniform convergence $p^{(\lambda)}_{\varepsilon}\to p_{\varepsilon}$ that the right-hand side converges to the same quantity for $X$, hence the bridges converge on any interval $[0,1-\varepsilon]$. Notice then that the bridges are invariant under space- and time-reversal (since Lévy processes are), so the part on the interval $[1-\varepsilon,1]$ converges as well, hence both are tight and the entire bridge on $[0,1]$ is tight. The convergence on any interval $[0,1-\varepsilon]$ then characterises uniquely the subsequential limits. Finally, the convergence of bridges for $\vartheta=0$ is transferred to the excursions by the Vervaat transform. ∎
In practice, the uniform convergence of transition densities can be often be checked by using the inverse Fourier transform. Precisely, we may write:
@@ 5.1.   Convergence of looptrees and labels
Let us first study the convergence of the looptrees and Gaussian labels associated with $X^{(\lambda)}$ and $X$. Recall from (2.6) and (2.8) respectively that they are defined using a parameter which acts as a multiplicative constant on the continuous part. In the next theorem, we fix the value of this constant: $1/2$ for looptrees and $1/3$ for the labels. We discuss this restriction after the statement.
@@
Recall that the factor $1/2$ in the looptree above refers to the multiplicative constant applied to the continuous part in (2.6) and similarly in (2.8) for $1/3$ in the label process. First notice that in absence of jumps in the limit, namely if $X$ is a scaled Brownian excursion, then for any $c>0$, the looptree $\mathscr{L}^{c}(X)=c\,\mathscr{L}^{1}(X)$ is a scaled Brownian tree and $Z^{c}=\sqrt{c}\,Z$ is a scaled Brownian snake so the choice of $c$ amounts to fix a normalisation. In the opposite case, if this continuous part in the limit vanishes, then the choice of this constant is irrelevant. Similarly, and this will be the case in our application, if the continuous part in $X^{(\lambda)}$ vanishes, then this constant is irrelevant in the sequence. The reason for this restriction is the following. The continuous part for $X$ appears as the limit of the continuous part of $X^{(\lambda)}$ together with sum of quantities of the form $R^{t}_{s}$ as defined in (2.3), corresponding to small jumps of $X^{(\lambda)}$. Now typically, the contribution to the looptree distances of these small jumps will be half of the sum of these $R^{t}_{s}$’s by symmetry, since the latter code the right-length of the cycles. Hence in order to control the looptrees with a different multiplicative constant than $1/2$, we must understand the precise contribution of the continuous part of $X^{(\lambda)}$ and its small jumps to the continuous part of $X$, whereas with the constant $1/2$, in the looptree, they both contribute equally.
@@
The last line converges in distribution to $d_{X}(s,s^{\prime})+d_{X}(t,t^{\prime})$ under the unconditional law or the bridge or excursion law by Lemma 5.1. This limit in turn is arbitrarily small when the pairs $(s,t)$ and $(s^{\prime},t^{\prime})$ are arbitrarily close to each other (for the Euclidean distance). Tightness of the looptree distances follows then easily from this control on the modulus of continuity.
Tightness of the label process is less immediate. Let us prove it providing that $d^{1/2}_{\mathscr{L}(X^{(\lambda)})}\to d^{1/2}_{\mathscr{L}(X)}$ in distribution, which we shall prove below. Our aim is again to control the modulus of continuity. Precisely, let us fix $\varepsilon,\theta>0$ and let us prove that there exists $\delta>0$ such that for every $\lambda$ large enough, we have:
@@
Let $C_{a}$ be the random variable as in (4.3) for the limit process $X$, for some fixed admissible $a>0$. Since $C_{a}<\infty$ almost surely then there exists a deterministic constant $M\in(0,\infty)$ such that $C_{a}\leq M$ with probability at least $1-\varepsilon$. Let $C>0$ which will be chosen large enough, depending only on $\theta$ and $M$ (and thus $\varepsilon$), and let us set $\kappa=(\theta/C)^{2}$ and then $N=\lceil(M/\kappa)^{1/a}\rceil=\lceil(MC^{2}/\theta^{2})^{1/a}\rceil$; we shall take $\delta=1/N$. Consider the event:
@@
($\mathop{}\!\mathbb{P}(\{\exists(s,t)\in[0,1]^{2}\colon|t-s|\leq 1/N\quad\text{and}\quad|Z^{(\lambda,1/3)}_{s}-Z^{(\lambda,1/3)}_{t}|>C\sqrt{\kappa}\}\cap G(N,\kappa,\lambda)).$
@@
$N\max_{0\leq i<N}\mathop{}\!\mathbb{P}(\{\exists s\in[0,1/N]\colon|Z^{(\lambda,1/3)}_{s+i/N}-Z^{(\lambda,1/3)}_{i/N}|>C\sqrt{\kappa}/4\}\cap G(N,\kappa,\lambda)).$
@@
$\leq K\int_{0}^{\sqrt{\kappa}}\sqrt{\log\lceil\sqrt{\kappa}/r\rceil}\mathop{}\!\mathrm{d}r$
$\leq K\sqrt{\kappa}\int_{0}^{1}\sqrt{\log\lceil 1/x\rceil}\mathop{}\!\mathrm{d}x,$
and the integral is some universal constant as well. Let us write $L\sqrt{\kappa}$ the last line. Recall the exponent $a<1/\boldsymbol{\eta}$ which we have fixed. If we were able to take $a>1/2$ then we could conclude by the Markov inequality. However in the case $\boldsymbol{\eta}=2$, in particular in the Brownian case, we need a slightly better bound. A very strong bound [BLM13, Theorem 5.8] which follows from standard Gaussian concentration results is the following: for every $u>0$, we have
$\mathop{}\!\mathbb{P}\Bigl{(}G(N,\kappa,\lambda)\cap\Bigl{\{}\sup_{s\in[0,1/N]}|Z^{(\lambda,1/3)}_{s+i/N}-Z^{(\lambda,1/3)}_{i/N}|>u+\mathop{}\!\mathbb{E}\Bigl{[}\sup_{s\in[0,1/N]}|Z^{(\lambda,1/3)}_{s+i/N}-Z^{(\lambda,1/3)}_{i/N}|\Bigm{|}X^{(\lambda)}\Bigr{]}\Bigr{\}}\Bigr{)}$
@@
$\mathop{}\!\mathbb{P}(\exists(s,t)\in[0,1]^{2}\colon|t-s|\leq 1/N\quad\text{and}\quad|Z^{(\lambda,1/3)}_{s}-Z^{(\lambda,1/3)}_{t}|>\theta)\leq 2\varepsilon+N\exp\Bigl{(}-\frac{(C/4-L)^{2}}{2}\Bigr{)}.$
@@
Proof of uniqueness of the limits for the unconditioned processes.
@@
Let us here consider the unconditioned Lévy processes and let us first prove that for any $t>0$ fixed, the looptree distance from 00 to $t$ for $X^{(\lambda)}$ converges in distribution to that for $X$. Then we discuss the convergence of the distance between any $s$ and $t$ fixed, the argument for more general finite dimensional distributions is the same. We then briefly mention the modifications to control the labels.
Step 1: distance from 00 to $t$. Fix $t>0$. Let us denote by $C^{(\lambda)}_{t}$ and $R^{(\lambda),t}_{s}$ respectively the continuous part and jump part that appear in the looptree distance (2.6) associated with $X^{(\lambda)}$. By standard properties of the Skorokhod $J_{1}$ topology, the convergence $X^{(\lambda)}\to X$ implies that for any $\delta>0$, the jumps larger than $\delta$ of $X$ are the limits of large jumps of $X^{(\lambda)}$, and quantities such that $R^{(\lambda),t}_{s}$ for such jump times $s$ converge as well to their analogue for $X$. Hence, the contribution of the times of jumps larger than $\delta$ in the looptree distance associated with $X^{(\lambda)}$ converges to that for $X$. It remains to consider the joint contribution of the continuous part and the small jumps, namely to prove that:
$\frac{1}{2}C^{(\lambda)}_{t}+\sum_{s\prec t}\min(R^{(\lambda),t}_{s},\Delta X^{(\lambda)}_{s}-R^{(\lambda),t}_{s})\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}\leq\delta\}}\enskip\mathop{\longrightarrow}^{(d)}\enskip\frac{1}{2}C_{t}$ (5.1)
as first $\lambda\to\infty$ and then $\delta\to 0$.
First note that the convergence in the Skorokhod $J_{1}$ topology also implies that $X^{(\lambda)}_{t-}$ converges to $X_{t-}$, hence as a consequence of the previous lines, we have for any $\delta>0$ fixed:

### v3

@@ 5.1.   Hölder continuity of the distances
Recall the moment bound on the Brownian labels (2.9), then if the Hölder continuity (5.1) on the looptree distances held true with a deterministic constant $C_{a}$, then we could apply the classical Kolmogorov continuity criterion, showing that for any $a<1/\boldsymbol{\eta}$ and any $q>0$, the process $(Z^{\mathrm{ex}}_{t})_{t}$ has a modification that is Hölder continuous with any exponent $b<q^{-1}(aq/2-1)$, which is arbitrarily close to $1/(2\boldsymbol{\eta})$ when $q$ is large and $a$ is close to $1/\boldsymbol{\eta}$.
In order to do so, let $A_{n}=\{n-1<C_{a}\leq n\}$ for each $n\geq 1$, so these events are disjoint and their union has probability $1$. Let us set =$\mathbb{N}^{\prime}=\{n\geq 1\colon\mathop{}\!\mathbb{P}(A_{n})\neq 0\}$ and $A_{0}=(\bigcup_{n\in\mathbb{N}^{\prime}}A_{n})^{c}$. Then for every $n\in\mathbb{N}^{\prime}$ we can apply Kolmogorov’s criterion under $\mathop{}\!\mathbb{P}(\,\cdot\mid A_{n})$ and deduce the existence of a modification $Z^{n}$ that is Hölder continuous with any exponent $b<1/(2\boldsymbol{\eta})$, under the conditional law. Let $Z^{\prime}$ coincide with $Z^{n}$ on $A_{n}$ for every $n\in\mathbb{N}^{\prime}$ and let $Z^{\prime}$ be a fixed arbitrary value on $A_{0}$. Then for every $t\in[0,1]$, since $\mathop{}\!\mathbb{P}(A_{0})=0$, then we have $\mathop{}\!\mathbb{P}(Z^{\mathrm{ex}}_{t}=Z^{\prime}_{t})=\sum_{n\in\mathbb{N}^{\prime}}\mathop{}\!\mathbb{P}(A_{n})\mathop{}\!\mathbb{P}(Z^{\mathrm{ex}}_{t}=Z^{n}\mid A_{n})=1$ so $Z^{\prime}$ is a modification of $Z^{\mathrm{ex}}$. By the same argument $Z^{\prime}$ is almost surely Hölder continuous with any exponent $b<1/(2\boldsymbol{\eta})$. ∎
From now on, we always work with these Hölder continuous modifications. This further implies that the distances in the map are Hölder continuous.
@@
For any $a\in(0,1/(2\boldsymbol{\eta}))$, there exists $C_{a}\in(0,\infty)$ such that:
@@
Similarly to Proposition 5.1, we shall see that the exponent $1/(2\boldsymbol{\eta})$ is optimal in the sense that almost surely, no such $C_{a}<\infty$ exist for $a>1/(2\boldsymbol{\eta})$, see again Remark 5.7.
@@
The claim follows from the upper bound $d_{M}\leq D^{\circ}$ from Theorem 2.3, where we recall that $D^{\circ}$ is defined by
$D^{\circ}(s,t)=D^{\circ}(t,s)=Z^{\mathrm{ex}}_{s}+Z^{\mathrm{ex}}_{t}-2\inf_{[s,t]}Z^{\mathrm{ex}}.$
Then the Hölder continuity of $Z^{\mathrm{ex}}$ from Corollary 5.5 transfers to $D^{\circ}$ which thus satisfies the upper bound (5.4), and thus so does $d_{M}$. ∎
@@ 6. Approximations in the continuum world
In this section we ask ourselves the following question. Suppose that for all $\lambda>0$ we are given a Lévy process $X^{(\lambda)}$ with no negative jump and infinite variation paths, with exponent $\psi^{(\lambda)}$, and another one $X$ with exponent $\psi$. Suppose that $X^{(\lambda)}$ and $X$ satisfy the integrability condition (1.3), so they all have continuous transition densities $p^{(\lambda)}_{t}(x)$ and $p_{t}(x)$ respectively, and we can define bridges and excursions as before. If these transition densities converge, then it is well known that the Lévy processes converge in law; but do the corresponding looptrees, Gaussian labels, and continuum maps converge as well? We prove the convergence of labelled looptrees in Section 6.1 and that of maps (along a subsequence) in Section 6.2.
One could for instance consider the case where $X^{(\lambda)}$ is an $\alpha$-stable spectrally positive Lévy process with drift $\vartheta$, as mentioned in Section 3, with $\alpha\in(1,2)$ and $\vartheta\in\mathbb{R}$ that may both vary. We will study this in a companion paper.
Let us say a few words on the techniques used to derive these results. If the limit $X$ has no Gaussian part, then it has the pure jump property from [Mar24] and Section 4 there answers positively our question, see alternatively [Kha22, Theorem 3 (iii)] for the convergence of looptrees in the pure jump case. However, here we consider the general setting and we need additional arguments. First, tightness of the looptree distances is easy but that of the labels is not: we argue by adapting the proof of the usual Kolmogorov criterion and using some general concentration bounds for Gaussian processes. We then turn to the finite-dimensional marginals, that is, the pairwise distances and joint labels of finitely many points in the looptree. The contribution of the large cycles can easily be controlled, and the contribution of the microscopic parts is proved by a sort of weak law of large numbers in the spirit of [CK14]. Indeed, informally, the Lévy excursion codes the length of paths in the looptree when one always follows the cycles on their right, whereas the true distance minimises left and right lengths. We show that for microscopic cycles, this basically divides their lengths by two. We actually argue on the unconditioned process first and then use a local absolute continuity argument for the excursion; the convergence of the densities is used for this transfer. Once the convergence of labelled looptrees is proved, the convergence of maps along a subsequence, and the identification of the Brownian sphere and tree in the extreme regimes, is routine as the usual arguments to control discrete maps can be adapted.
@@
Suppose that for every $t\geq 0$, the transition density $p^{(\lambda)}_{t}$ converges uniformly to $p_{t}$ as $\lambda\to\infty$. Then $X^{(\lambda)}\to X$ in distribution for the Skorokhod $J_{1}$ topology under the unconditional law as well as under $\mathop{}\!\mathbb{P}(\,\cdot\mid X^{(\lambda)}_{1}=\vartheta)$ for any $\vartheta\in\mathbb{R}$, and finally under the excursion law.
@@
for any continuous and bounded function $F$. We infer from the convergence in distribution of the unconditioned processes and the uniform convergence $p^{(\lambda)}_{\varepsilon}\to p_{\varepsilon}$ that the right-hand side converges to the same quantity for $X$, hence the bridges converge on any interval $[0,1-\varepsilon]$. Notice then that the bridges are invariant under space- and time-reversal (since Lévy processes are), so the part on the interval $[1-\varepsilon,1]$ converges as well, and in particular it is tight and so the entire bridge on $[0,1]$ is tight. The convergence on any interval $[0,1-\varepsilon]$ then characterises uniquely the subsequential limits. Finally, the convergence of bridges for $\vartheta=0$ is transferred to the excursions by the Vervaat transform. ∎
In practice, the uniform convergence of transition densities can often be checked by using the inverse Fourier transform. Precisely, we may write:
@@ 6.1.   Convergence of looptrees and labels
Let us first study the convergence of the looptrees and Gaussian labels associated with $X^{(\lambda)}$ and $X$. Recall from (2.6) and (2.8) respectively that they are defined using a parameter that acts as a multiplicative constant on the continuous part. In the next theorem, we fix the value of this constant: $1/2$ for looptrees and $1/3$ for the labels. We discuss this restriction after the statement.
@@
As was previously mentioned, in the particular case where $X$ has no Gaussian part, then it has the pure jump property and the convergence of looptrees follows from [Mar24, Section 4], or alternatively [Kha22, Theorem 1.4]. In the general case, Theorem 1.6 in [Kha22] proves the convergence of a ‘shuffled’ version of the looptrees, provided that the excursions converge and under some technical assumption on this shuffling. This technical assumption is proved to hold for stable looptrees, and also that their shuffled version has the same law as the original looptree [Kha22, Theorem 6.6]. A possible alternative proof of the convergence in distribution of looptrees in Theorem 6.2 could thus be to extend this approach to Lévy looptrees, as noted in [Kha22, Remark 6.8]. This approach, however, does not seem to directly give joint convergence of the looptrees and the Lévy processes.
Let us discuss the choice of constants in this theorem. First notice that in absence of jumps in the limit, namely if $X$ is a scaled Brownian path, then for any $c>0$, the looptree $\mathscr{L}^{c}(X)=c\,\mathscr{L}^{1}(X)$ is a scaled Brownian tree and $Z^{c}=\sqrt{c}\,Z$ is a scaled Brownian snake so the choice of $c$ amounts to fix a normalisation. In the opposite case, if this continuous part in the limit vanishes, then the choice of this constant is irrelevant. Similarly, if the continuous part in $X^{(\lambda)}$ vanishes, then this constant is irrelevant in the sequence. The reason for this restriction is the following. The continuous part for $X$ appears as the limit of the continuous part of $X^{(\lambda)}$ together with sum of quantities of the form $R^{t}_{s}$ as defined in (2.3), corresponding to small jumps of $X^{(\lambda)}$. Now typically, the contribution to the looptree distances of these small jumps will be half of the sum of these $R^{t}_{s}$’s by symmetry, since the latter code the right-length of the cycles. Hence in order to control the looptrees with a different multiplicative constant than $1/2$, we must understand the precise contribution of the continuous part of $X^{(\lambda)}$ and its small jumps to the continuous part of $X$, whereas with the constant $1/2$, in the looptree, they both contribute equally. The same applies for the constant $1/3$ for the Gaussian labels.
@@
The last line converges in distribution to $d_{X}(s,s^{\prime})+d_{X}(t,t^{\prime})$ under the unconditional law or the bridge or excursion law by Lemma 6.1. This limit in turn is arbitrarily small when the pairs $(s,t)$ and $(s^{\prime},t^{\prime})$ are arbitrarily close to each other (for the Euclidean distance). Tightness of the looptree distances then follows easily from this control on the modulus of continuity.
Tightness of the label process is less immediate. Let us prove it provided that $d^{1/2}_{\mathscr{L}(X^{(\lambda)})}\to d^{1/2}_{\mathscr{L}(X)}$ in distribution, which we shall prove below. Our aim is again to control the modulus of continuity. Precisely, let us fix $\varepsilon,\theta>0$ and let us prove that there exists $\delta>0$ such that for every $\lambda$ large enough, we have:
@@
Let $C_{a}$ be the random variable as in (5.1) for the limit process $X$, for some fixed admissible $a>0$. Since $C_{a}<\infty$ almost surely then there exists a deterministic constant $M\in(0,\infty)$ such that $C_{a}\leq M$ with probability at least $1-\varepsilon$. Let $C>0$ which will be chosen large enough, depending only on $\theta$ and $M$ (and thus $\varepsilon$), and let us set $\kappa=(\theta/C)^{2}$ and then $N=\lceil(M/\kappa)^{1/a}\rceil=\lceil(MC^{2}/\theta^{2})^{1/a}\rceil$; we shall take $\delta=1/N$. Consider the good event:
@@
$\mathop{}\!\mathbb{P}(G(N,\kappa,\lambda)\cap\{\exists(s,t)\in[0,1]^{2}\colon|t-s|\leq 1/N\enskip\text{and}\enskip|Z^{(\lambda,1/3)}_{s}-Z^{(\lambda,1/3)}_{t}|>C\sqrt{\kappa}\}).$
@@
$N\max_{0\leq i<N}\mathop{}\!\mathbb{P}(G(N,\kappa,\lambda)\cap\{\exists s\in[0,1/N]\colon|Z^{(\lambda,1/3)}_{s+i/N}-Z^{(\lambda,1/3)}_{i/N}|>C\sqrt{\kappa}/4\}).$
@@
[…]
$\qquad\leq K\int_{0}^{\sqrt{\kappa}}\sqrt{\log\lceil\sqrt{\kappa}/r\rceil}\mathop{}\!\mathrm{d}r\leq K\sqrt{\kappa}\int_{0}^{1}\sqrt{\log\lceil 1/x\rceil}\mathop{}\!\mathrm{d}x,$
and the integral is some universal constant as well. Let us write $L\sqrt{\kappa}$ this upper bound. Recall the exponent $a<1/\boldsymbol{\eta}$ which we have fixed. If we were able to take $a>1/2$ then we could conclude by the Markov inequality. However in the case $\boldsymbol{\eta}=2$, in particular in the Brownian case, we need a slightly better bound. A very strong bound [BLM13, Theorem 5.8] which follows from standard Gaussian concentration results is the following: for every $u>0$, we have
$\mathop{}\!\mathbb{P}\Bigl{(}G(N,\kappa,\lambda)\cap\Bigl{\{}\sup_{s\in[0,\frac{1}{N}]}|Z^{(\lambda,\frac{1}{3})}_{s+i/N}-Z^{(\lambda,\frac{1}{3})}_{i/N}|>u+\mathop{}\!\mathbb{E}\Bigl{[}\sup_{s\in[0,\frac{1}{N}]}|Z^{(\lambda,\frac{1}{3})}_{s+i/N}-Z^{(\lambda,\frac{1}{3})}_{i/N}|\Bigm{|}X^{(\lambda)}\Bigr{]}\Bigr{\}}\Bigr{)}$
@@
$\mathbb{P}\mathopen{}\mathclose{{}\left(\exists(s,t)\in[0,1]^{2}\colon|t-s|\leq\frac{1}{N}\,\text{ and }\,|Z^{(\lambda,\frac{1}{3})}_{s}-Z^{(\lambda,\frac{1}{3})}_{t}|>\theta}\right)\leq 2\varepsilon+N\exp\Bigl{(}-\frac{(\frac{C}{4}-L)^{2}}{2}\Bigr{)}.$
@@
Proof of convergence for unconditioned processes.
@@
Let us first consider the unconditioned Lévy processes and let us prove that for any $t>0$ fixed, the looptree distance from 00 to $t$ for $X^{(\lambda)}$ converges in distribution to that for $X$. Then we discuss the convergence of the distance between any $s$ and $t$ fixed, the argument for more general finite dimensional distributions is the same. We then briefly mention the modifications to control the labels.
Step 1: distance from 00 to $t$. The distance from 00 to $t$ in the looptree coded by $X$ equals
$\frac{1}{2}C_{t}+\sum_{s\prec t}\min(R^{t}_{s},\Delta X_{s}-R^{t}_{s}).$
Let us denote by $C^{(\lambda)}_{t}$ and $R^{(\lambda),t}_{s}$ respectively the continuous part and jump part that appear similarly in the looptree distance associated with $X^{(\lambda)}$. We shall argue that, when $\delta>0$ is small, the following terms
$\sum_{s\prec t}\min(R^{(\lambda),t}_{s},\Delta X^{(\lambda)}_{s}-R^{(\lambda),t}_{s})\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}\leq\delta\}}\qquad\text{and}\qquad\frac{1}{2}\sum_{s\prec t}R^{(\lambda),t}_{s}\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}\leq\delta\}}$
are close to each other, and that the second one will contribute, together with $\frac{1}{2}C^{(\lambda)}_{t}$, to the continuous part $\frac{1}{2}C_{t}$ in the limit as $\lambda\to\infty$ and $\delta\to 0$. We shall implicitly restrict ourselves to values of $\delta$ which are not atoms of the Lévy measure $\pi$. This set of atoms is at most countable, so we can indeed find a sequence outside of it that tends to 00. Henceforth, we thus assume that $X$ has no jump of size exactly $\delta$.
Recall that $C_{t}=X_{t-}-\sum_{s\prec t}R^{t}_{s}$, and the same holds for $X^{(\lambda)}$. Standard properties of the Skorokhod $J_{1}$ topology imply that, since $X^{(\lambda)}$ converges in distribution to $X$, then $X^{(\lambda)}_{t-}\to X_{t-}$ and, jointly, the (finitely many) jumps larger than $\delta$ of $X$ all arise as limits of jumps larger than $\delta$ of $X^{(\lambda)}$ (without boundary effect since $X$ has no jump of size exactly $\delta$), and the times at which they take place converge, and so do the associated quantities $R^{(\lambda),t}_{s}$. As a consequence,
$C^{(\lambda)}_{t}+\sum_{s\prec t}R^{(\lambda),t}_{s}\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}\leq\delta\}}$ $=$ $X^{(\lambda)}_{t-}-\sum_{s\prec t}R^{(\lambda),t}_{s}\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}>\delta\}}$